OpenStudy (anonymous):

How does √3 √3 equal 3 since both square roots are undefined?

jimthompson5910 (jim_thompson5910):

how are they undefined?

OpenStudy (anonymous):

I thought √3 is undefined

OpenStudy (anonymous):

and so is the other one

jimthompson5910 (jim_thompson5910):

the square root of 3 is some number (roughly 1.7) so it's definitely defined

OpenStudy (anonymous):

Oh okay

OpenStudy (anonymous):

But how do they equal 3?

jimthompson5910 (jim_thompson5910):

\[\Large \sqrt{3}*\sqrt{3} = (\sqrt{3})^2 = 3\] the square and square root cancel each other out

OpenStudy (anonymous):

Aaaah I see

jimthompson5910 (jim_thompson5910):

Another way \[\Large \sqrt{3}*\sqrt{3} = \sqrt{3*3} = \sqrt{9} = 3\]

OpenStudy (anonymous):

thanks for the help :)

jimthompson5910 (jim_thompson5910):

np